OpenStudy (annjewell1998):

The area of a trapezoid is A=0.5(b1+b2)h find b1 if A=100, h=5, and b2=12. Please help.

Directrix (directrix):

A = 0.5 * (b1 + b2 ) *h find b1 if A=100, h=5, and b2=12.

Directrix (directrix):

100 = .5 * (b1 + 12) * 5 100 = .5 * (b1 + 12) *5 Multiply both sides by 2 200 = 2* .5 * (b1 + 12) *5 200 = 1 * (b1 + 12)*5 200 = (b1 + 12) * 5 Divide both sides by 5 40 = b1 + 12 b1 = ? @annjewell1998

OpenStudy (annjewell1998):

I have no idea. do i divide 40 by 5 and 1 by 5? i tried and it didn't look right.

Directrix (directrix):

You should read the entire thread and follow the steps rather than looking at the last post.

Directrix (directrix):

40 = b1 + 12 b1 = ? @annjewell1998 What is 40 minus 12 = ? That is b1.